OpenStudy (anonymous):

Solve and graph. 2|-2x+7|>=18 Would it be x<=-8 and x>=-1???

OpenStudy (mertsj):

\[|-2x+7|\ge9\] \[-2x+7\ge9 or -2x+7\le-9\]

OpenStudy (mertsj):

Solve both parts.

jimthompson5910 (jim_thompson5910):

2|-2x+7|>=18 |-2x+7|>=18/2 |-2x+7|>=9 -2x+7>=9 or -2x+7<=-9 -2x>=9-7 or -2x<=-9-7 -2x>=2 or -2x<=-16 x<=2/(-2) or x>=-16/(-2) <<<--- this is probably where you made your mistake x <= -1 or x >= 8